zackariahharriett:

4(2x - 5) + 15 = 11 Write the steps you will use to solve the equation and explain each step

smackz:

ok so first thing yo uwant to do is use pemdas process.

smackz:

so first we di=eal whjats in parantheses

smackz:

so we will distirbute the 4 among the positive 2x, and the negative 5 which becomes 8x - 5 + 15

smackz:

so now u wanan combine like terms, which would be 15+ -5 = positive 10

smackz:

so now u have 8x + 10 = 11

smackz:

now u want to a lil subtracting which leaves u with 8x + 10 = 11 - 10 - 10 because we want the x to be by itself

smackz:

that gives u 8x = 1 do dividing the answer yo uget is 8 because 8 can divide itself one time ;) hope this helps